You are to find one of two numbers that is two times the size of the first number. These two numbers together equal 180. What is that number?

Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

You are to find one of two numbers that is two times the size of the first number. These two numbers together equal 180. What is that number?
Equation:
a + b = 180
b = 2a
----
Substitute and solve for "a":
a + 2a = 180
3a = 60
a = 20
b = 160
============
Cheers,
Stan H.
============

Answer by nerdybill(7384)

You are to find one of two numbers that is two times the size of the first number. These two numbers together equal 180. What is that number?
.
Let x = first number
then
2x = second number
.
from:"two numbers together equal 180"
x + 2x = 180
3x = 180
x = 180/3
x = 60 (answer)